how to draw relation between velocity and acceleration when velocity is a function of displacement
 

Grade:Select Grade

1 Answers

Nicho priyatham
8 years ago
a=dv/dt
so a=(dv/ds)(ds/dt)  so v=ds/dt
a=v(dv/ds)
use the above equation 2 get the relation